# Kink Finder at Belle II

### v1

@@ 5.1 Reconstructed track pairs
The main parameters of the in-flight-decay kink are the decay vertex position (x,y,z)(x,\,y,\,z), the corresponding momenta of the mother and the daughter particles, and the momentum of the daughter particle in the mother particle rest frame. Figure 2 shows the vertex $r$- and $z$-component resolutions for kaon and pion decays. The overall resolution of the vertex position, defined as the spread of the distribution of the distance from the real decay vertex to the reconstructed one, is 0.9 cm for kaon decays and 2.4 cm for pion decays. Approximately 93% and 62% of events are contained within a 1 cm region in the $rz$-plane, respectively. A better resolution for kaons is attributed to the larger energy release in the decay, leading to a more prominent kink that is easier to fit.
Figure 2: Resolution of the $r$- and $z$-components of the vertex position. Panel (a, b) are obtained with the $\tau^{-}\rightarrow K^{-}\nu_{\tau}$ sample and panel (c, d) with the $\tau^{-}\rightarrow\pi^{-}\nu_{\tau}$ sample.
The daughter momentum resolutions in the laboratory frame ($p^{\text{lab}}$) and in the mother particle rest frame ($p_{dm}$) for kaon and pion two-body decays are shown in Fig. 3 and 4, respectively. For both samples, a significant improvement is obtained with the application of the Kink Finder: better resolutions and bias elimination caused by energy losses correction for the default case. Figure 3(c) shows two distinct peaks originating from $K^{-}\rightarrow\pi^{-}\pi^{0}$ and $K^{-}\rightarrow\mu^{-}\bar{\nu}_{\mu}$ decays. Similarly, a single peak from $\pi^{-}\rightarrow\mu^{-}\bar{\nu}_{\mu}$ decay is clearly identified in Fig. 4(c).
@@
In Fig. 5, we illustrate the composition of the reconstructed kinks in the $B\bar{B}$ sample by providing the distributions of the daughter particle momentum in the mother particle rest frame for two pairs of different mass hypotheses: muon (pion) and pion (kaon) for daughter (mother) particles, corresponding to $\pi^{-}\rightarrow\mu^{-}\bar{\nu}_{\mu}$ and $K^{-}\rightarrow\pi^{-}\pi^{0}$ decays. Here, we have further suppressed the contribution from clones, using the output of the combined fit described in Sec. 4.2. Narrow peaks from two-body pion and kaon decays are clearly discerned. These distributions can be compared with the corresponding ones from Belle, obtained without the Kink Finder [Belle:2023dyc]. A significant improvement in the resolution is observed, which creates better conditions for physics studies, e.g., the Michel parameter $\xi^{\prime}$ measurement.
Figure 5: Daughter particle momentum in the mother rest frame with muon and pion mass hypotheses (a) and pion and kaon mass hypotheses (b). The various sources of kinks in $B\bar{B}$ sample are shown as stacked histograms.
@@ 5.2 Combined tracks
In addition, the mere fact that a track has been split by the Kink Finder is sufficient to reduce the PID fake rate. The fake rate of the standard Belle II Neural-Network-based binary PID [Belle-II:2025tpe] is studied for kaons and pions using the $B\bar{B}$ sample. The exclusion of particles found by the Kink Finder improves the overall kaon and pion fake rates by 1% and 5%, respectively; however, the effect depends on the transverse momentum $p_{T}$. For pions, the main improvement is in the low-$p_{T}$ region $p_{T}<1.2\,\text{GeV}/c$, with a maximum of 10% for $p_{T}<0.6\,\text{GeV}/c$. For kaons, the improvement is negligible for the interval $0.6\,\text{GeV}/c<p_{T}<1.2\,\text{GeV}/c$, while for the remaining region, it is around 2–3%. We have also studied the potential maximal reduction of the fake rate if all in-flight decays, reconstructed as one combined track, are found by the Kink Finder. For pions, the largest effect is again obtained for low $p_{T}$ particles with a maximum of 15% for $p_{T}<0.6\,\text{GeV}/c$. For kaons, the expected improvement is $\approx 3\%$ in the same region as for the veto on tracks found by the Kink Finder. Thus, we conclude that the in-flight decays in general have quite a small effect on the PID fake rate.
@@ 6 Problems and further improvement
While the mother track is reconstructed in the majority of kink events, daughter tracks have relatively low efficiency. Studies show that the application of the Belle II segment finder and the following local track-finder [Bertacchi:2020eez] can increase the daughter track reconstruction efficiency by up to a factor of two. However, at present, the local track-finder is switched off due to high fake rate and clone rate. An upgrade, including the re-enabling of this feature, is under discussion.
The Graph Neural Network (GNN) track-finding, recently developed and studied at Belle II [Reuter:2024kja], has shown an impressive improvement in the finding efficiency of the secondary tracks. Presently, this algorithm is being implemented in basf2. It is also considered to integrate kinks in the GNN training to increase the Kink Finder efficiency.
@@
Finally, the precision of the kink parameters can be enhanced by using the geometric constraint of coupled helices at the vertex. A Kalman filter, applied to all hits using the information about the continuity of the kink (inspired by Ref. [Astier:1999rs]), would improve vertex and momentum resolutions. Implementation within the genfit2 package requires further development and is left for future updates of the Kink Finder.
@@ 7 Conclusion
In conclusion, the first version of the Kink Finder has been developed and implemented in the Belle II tracking software. We have provided a detailed description of the algorithms and conducted a performance study using MC events. The overall efficiency of the reconstruction of the in-flight-decay kinks is obtained to be around $40\%$ for both kaon and pion decays. The resolution of the kink track parameters is improved compared to the default case, where the kink can be reconstructed at the analysis level without the Kink Finder. The obtained efficiency and resolutions can significantly improve the precision of the Michel parameter $\xi^{\prime}$ measurement compared to the current result by Belle. The number of cloned tracks identified by Kink Finder is reduced by a factor of two. The PID fake rate is significantly reduced for low-energy pions while the corresponding impact in kaons is marginal. We have identified several ways to further improve the Kink Finder and will implement these in the near future.
@@ Acknowledgements
The work is supported by the National Natural Science Foundation of China (NSFC) under Contracts Nos. 12335004 and 12575092. The study has been funded within the framework of the HSE University Basic Research Program. The article was prepared within the framework of the project “Mirror Laboratories” HSE University.

### v2

@@ 5.1 Reconstructed track pairs
The main parameters of the in-flight-decay kink are the decay vertex position (x,y,z)(x,\,y,\,z), the corresponding momenta of the mother and the daughter particles, and the momentum of the daughter particle in the mother particle rest frame. Figure 2 shows the kink vertex $r$- and $z$-component resolutions for kaon and pion decays. The overall resolution of the vertex position, defined as the spread of the distribution of the distance from the real decay vertex to the reconstructed one, is 0.9 cm for kaon decays and 2.4 cm for pion decays. Approximately 93% and 62% of events are contained within a 1 cm region in the $rz$-plane, respectively. A better resolution for kaons is attributed to the larger energy release in the decay, leading to a more prominent kink that is easier to fit.
Figure 2: Resolution of the $r$- and $z$-components of the kink vertex position for in-flight decays. Panel (a, b) are obtained with the $\tau^{-}\rightarrow K^{-}\nu_{\tau}$ sample and panel (c, d) with the $\tau^{-}\rightarrow\pi^{-}\nu_{\tau}$ sample.
The daughter momentum resolutions in the laboratory frame ($p^{\text{lab}}$) and in the mother particle rest frame ($p_{dm}$) for kaon and pion two-body decays are shown in Fig. 3 and 4, respectively.
@@
For both samples, a significant improvement is obtained with the application of the Kink Finder. Better resolutions for the daughter track parameters are achieved with the hit reassignment and improved estimation of the track seed time used for the track extrapolation based on the information from the last hit of the mother track. For the daughter particle momentum in the mother particle rest frame, there is an additional positive effect from the better accuracy of the decay vertex measurement, improving the determination of the angle between daughter and mother momenta. Finally, in the default case, there is a bias of $2{\mathrm{\,Me\kern-1.00006ptV\!/}c}$ ($3{\mathrm{\,Me\kern-1.00006ptV\!/}c}$) in the transverse momentum of the daughter particle and $9{\mathrm{\,Me\kern-1.00006ptV\!/}c}$ ($2{\mathrm{\,Me\kern-1.00006ptV\!/}c}$) in the $p_{dm}$ momentum for kaon (pion) decays caused by energy losses corrections. This bias is completely (up to statistical uncertainty) eliminated by the Kink Finder. Figure 3(c) shows two distinct peaks originating from $K^{-}\rightarrow\pi^{-}\pi^{0}$ and $K^{-}\rightarrow\mu^{-}\bar{\nu}_{\mu}$ decays. Similarly, a single peak from $\pi^{-}\rightarrow\mu^{-}\bar{\nu}_{\mu}$ decay is clearly identified in Fig. 4(c).
@@
In Fig. 5, we illustrate the composition of the reconstructed kinks in the $B\bar{B}$ sample by providing the distributions of the daughter particle momentum in the mother particle rest frame for two pairs of different mass hypotheses: muon (pion) and pion (kaon) for daughter (mother) particles, corresponding to $\pi^{-}\rightarrow\mu^{-}\bar{\nu}_{\mu}$ and $K^{-}\rightarrow\pi^{-}\pi^{0}$ decays.
Figure 5: Daughter particle momentum in the mother’s rest frame with muon and pion mass hypotheses (a) and pion and kaon mass hypotheses (b). The various sources of kinks in $B\bar{B}$ sample are shown as stacked histograms.
Here, we have further suppressed the contribution from clones, using the output of the combined fit described in Sec. 4.1. Narrow peaks from two-body pion and kaon decays are clearly discerned. These distributions can be compared with the corresponding ones from Belle, obtained without the Kink Finder [15]. A significant improvement in the resolution is observed, which creates better conditions for physics studies, e.g., the Michel parameter $\xi^{\prime}$ measurement.
@@ 5.2 Combined tracks
In addition, the mere fact that a track has been split by the Kink Finder is sufficient to reduce the PID fake rate. The fake rate of the standard Belle II Neural-Network-based binary PID [4] is studied for kaons and pions using the $B\bar{B}$ sample. The exclusion of particles found by the Kink Finder improves the overall kaon and pion fake rates by 1% and 5%, respectively; however, the effect depends on the transverse momentum $p_{T}$. For pions, the main improvement is in the low-$p_{T}$ region $p_{T}<1.2\,\text{GeV}/c$, with a maximum of 10% for $p_{T}<0.6\,\text{GeV}/c$. For kaons, the improvement is negligible for the interval $0.6\,\text{GeV}/c<p_{T}<1.2\,\text{GeV}/c$, while for the remaining region, it is around 2–3%. We have also studied the potential maximal reduction of the fake rate if all in-flight decays, reconstructed as one combined track, are found by the Kink Finder. For pions, the largest effect is again obtained for low $p_{T}$ particles with a maximum of 15% for $p_{T}<0.6\,\text{GeV}/c$. For kaons, the expected improvement is $\approx 3\%$ in the same region as for the veto on tracks found by the Kink Finder. Thus, we conclude that the in-flight decays in general have quite a small effect on the PID fake rate.
@@ 6 Problems and further improvement
While the mother track is reconstructed in the majority of kink events, daughter tracks have relatively low efficiency. Studies show that the application of the Belle II segment finder and the following local track-finder [12] can increase the daughter track reconstruction efficiency by up to a factor of two. However, at present, the local track-finder is switched off due to high fake rate and clone rate. An upgrade, including the re-enabling of this feature, is under discussion.
The Graph Neural Network (GNN) track-finding, recently developed and studied at Belle II [29], has shown an impressive improvement in the finding efficiency of the secondary tracks. Presently, this algorithm is being implemented in basf2. It is also considered to integrate kinks in the GNN training to increase the Kink Finder efficiency.
@@
Finally, the precision of the kink parameters can be enhanced by using the geometric constraint of coupled helices at the vertex. A Kalman filter, applied to all hits using the information about the continuity of the kink (inspired by Ref. [10]), would improve vertex and momentum resolutions. Implementation of two features described above within the genfit2 package requires further development and is left for future updates of the Kink Finder.
@@ 7 Conclusion
In conclusion, the first version of the Kink Finder has been developed and implemented in the Belle II tracking software. We have provided a detailed description of the algorithms and conducted a performance study using MC events. The overall efficiency of the reconstruction of the in-flight-decay kinks is obtained to be around $40\%$ for both kaon and pion decays. The resolution of the kink track parameters is improved compared to the default case, where the kink can be reconstructed at the analysis level without the Kink Finder. The obtained efficiency and resolutions can significantly improve the precision of the Michel parameter $\xi^{\prime}$ measurement compared to the current result by Belle. The number of cloned tracks identified by Kink Finder is reduced by a factor of two. The PID fake rate is reduced for low-energy pions while the corresponding impact in kaons is marginal. We have identified several ways to further improve the Kink Finder and will implement these in the near future.
Regarding the application of the Kink Finder for the data analysis, the Belle II data has not been reprocessed with the Kink Finder yet due to the Belle II software release schedule. According to the current plans, it may take at least 1 or 2 years from now until the data with the applied Kink Finder is available for analysts. Even more time will be required before the first analyses of kinks are complete.
@@ Acknowledgements
The work is supported by the National Natural Science Foundation of China (NSFC) under Contracts Nos. 12335004 and 12575092. The article was prepared within the framework of the project “Mirror Laboratories” HSE University.